Solve the equation.
2p^3 - 7p^2 - 50p + 175 = 0
I tried factoring out a p and then FOILing the remaining part while ignoring the " + 175 " part, but it didn't work. :(

Answer by jim_thompson5910(35256)

2p^3 - 7p^2 - 50p + 175 = 0


(2p^3 - 7p^2) + (-50p + 175) = 0


p^2(2p - 7) -25 (2p - 7) = 0


(p^2-25)(2p-7) = 0


(p-5)(p+5)(2p-7) = 0


I'll let you take over from here.